# Byzantine-Robust Gossip: Insights from a Dual Approach

### v1

@@ Abstract
Distributed approaches have many computational benefits, but they are vulnerable to attacks from a subset of devices transmitting incorrect information. This paper investigates Byzantine-resilient algorithms in a decentralized setting, where devices communicate directly with one another. We leverage the so-called dual approach to design a general robust decentralized optimization method. We provide both global and local clipping rules in the special case of average consensus, with tight convergence guarantees. These clipping rules are practical, and yield results that finely characterize the impact of Byzantine nodes, highlighting for instance a qualitative difference in convergence between global and local clipping thresholds. Lastly, we demonstrate that they can serve as a basis for designing efficient attacks.
Machine Learning, Decentralized optimization, Byzantine Robustness, ICML
@@ 1 Introduction
As datasets become larger and more widely distributed across computing devices, distributed optimization becomes increasingly important. Decentralized optimization offers a solution to leverage the distributed computing power without suffering from communication bottlenecks or privacy leaks induced by a central server. However, communicating with large numbers of low-confidence devices opens up new security problems. Optimization can be thwarted by devices that send malicious messages, either deliberately because they are controlled by hostile parties, or because software and hardware problems cause them to behave incorrectly. We consider the problem of optimizing a finite sum of local functions distributed over the nodes of a network, where we assume some nodes to be Byzantine: instead of properly executing the algorithm, they modify their messages to disturb the optimization procedure. Robustness to Byzantine attacks is a worst-case scenario in the sense that we model adversarial nodes as omniscient, able to collude, and capable of sending different information to different nodes (Lamport et al., 1982; Blanchard et al., 2017)
We assume that $n$ nodes are linked by a communication network which we represent as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with $\mathcal{V}=\{1,\dots,n\}$ the vertices and $\mathcal{E}$ the edges. In the Byzantine context, nodes (i.e. vertices) are split into honest nodes $\mathcal{V}_{h}$ and Byzantine ones $\mathcal{V}_{b}$. Note that since Byzantine nodes are assumed to be omniscient and can send any values, edges between Byzantine nodes do not matter: they do not provide them additional information nor allow them to send different messages. Therefore, we only consider honest-to-honest edges $\mathcal{E}_{h}$, and honest-to-Byzantine ones $\mathcal{E}_{b}$. We always assume the subgraph of honest nodes $\mathcal{G}_{h}:=(\mathcal{V}_{h},\mathcal{E}_{h})$ to be connected. For any $i\in\mathcal{V}_{h}$, we denote $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ the local objective function on node $i$. Then, the optimization problem writes:
@@
We solve this problem in a decentralized way. Each node $i\in\mathcal{V}$ maintains and updates a local parameter $x_{i}\in\mathbb{R}^{d}$. We consider the standard gossip approaches (Boyd et al., 2006; Nedic & Ozdaglar, 2009; Duchi et al., 2011; Scaman et al., 2017), in which communication steps consist in performing local parameter averages between neighbors. This corresponds to applying a so-called gossip matrix that encapsulates the topology of the graph.
@@
Such gossip algorithms naturally arise when solving a well-chosen dual formulation of Equation 1 with standard first-order methods. This dual approach gives a principled framework to design and analyze efficient decentralized algorithms, successfully enabling for instance acceleration (Scaman et al., 2017; Kovalev et al., 2020) or variance reduction (Hendrikx et al., 2019). In this paper, we investigate the benefits of this approach in the Byzantine robust setting.
Since its introduction by Blanchard et al. (2017), the setting of Byzantine-robust learning has been extensively studied. Yet,, most works focus on the federated case, in which all nodes directly communicate with a central server (Yin et al., 2018; Chen et al., 2017; Alistarh et al., 2018; Li et al., 2019; El-Mhamdi et al., 2020; Karimireddy et al., 2020, 2021; Farhadkhani et al., 2022), or in the fully connected case, in which any pair of agents may communicate (El-Mhamdi et al., 2021; Farhadkhani et al., 2023). These approaches typically rely on federated stochastic gradient methods, coupled with robust aggregation. In the federated case, specific attacks have been proposed by leveraging the stochasticity of the problem (Baruch et al., 2019; Xie et al., 2020).
While the centralized case has been studied extensively, the fully decentralized setting has received less attention. A keystone to Byzantine-robust decentralized learning is to find a robust approximate agreement, which has been studied in 2012-2014 (Vaidya et al., 2012; Vaidya, 2014; LeBlanc et al., 2013). Yet these works only focused on achieving agreement within the convex hull of the initial parameters, without seeking optimality. A few recent works have managed to propose Byzantine-robust decentralized learning algorithms. Such methods typically rely on adapting federated robust aggregation schemes (Peng et al., 2021; Fang et al., 2022). More recently, aggregation rules leveraging the difference of neighbors positions have been successfully proposed (He et al., 2022; Wu et al., 2023; Farhadkhani et al., 2023). We summarize the properties of these works in Table 1. However, despite sufficient robustness criteria for adapted rules (Wu et al., 2023; Kuwaranancharoen & Sundaram, 2023) and impossibility results (El-Mhamdi et al., 2021; He et al., 2022), a fine characterization of the influence of Byzantine nodes remains elusive, even for basic problems such as gossip averaging.
Practical Sparse $\mathcal{G}$ Convergence Dual
Farhadkhani et al. (2023)
He et al. (2022)
Wu et al. (2023)
Peng et al. (2021)
ECG - Global-Clipping
ECG -  Local-Clipping
Table 1: Comparison to main references. ECG holds for $\operatorname{\operatorname{EdgeClippedGossip}}$, i.e., Algorithm 1. Practical: indicates that the algorithmic rule can be used without oracle knowledge. Sparse $\mathcal{G}$: indicates that the methods can be used for any graph. Convergence: indicates that convergence is always ensured up to a neighborhood, generally of size $\delta_{\max}\zeta^{2}$. Dual: indicates that the algorithm is grounded in the dual approach.
Contributions. In this paper, (i) we demonstrate how the dual approach can be used to interpret clipped Gossip, an existing robust algorithm, as a clipped dual gradient descent, (ii) derive clipping rules for this algorithm, (iii) provide convergence guarantees to better understand their robustness (or there lack of) to Byzantine attacks. In particular, we highlight two main settings depending on whether we choose the clipping threshold locally or globally. We obtain non-asymptotic convergence results for both versions, and discuss pros and cons of each method. Note that that our clipping rules can be implemented in practice (assuming some coordination for the global one). (iv) We eventually propose a constructive approach to design attacks by exploiting the topology of the network.
Our main algorithm, $\operatorname{\operatorname{EdgeClippedGossip}}$, is defined for generic objective functions. Yet, we restrict its analysis to the problem of decentralized mean-estimation, which allows us to obtain tight convergence bounds. We insist on the fact that this seemingly simple problem is the base building block for all decentralized algorithms, and yet is far from being understood in the Byzantine setting. The general (convex) case is left as an important open problem, though the dual approach allows to efficiently disentangle communication and computation in decentralized optimization (Hendrikx et al., 2019; Kovalev et al., 2020), and is thus a promising approach for a next step.
@@
2 ClippedGossip and its dual interpretation
This section introduces a dual point of view on clipped gossip. We start by introducing the basics of gossip algorithms and their flaws in the Byzantine setting. Then, we present the algorithm proposed by He et al. (2022). Finally we derive our algorithm of interest, $\operatorname{\operatorname{EdgeClippedGossip}}$, which in a particular case recovers their algorithm enabling to reinterpret it as Clipped Gradient Descent in the dual.
Setting: We denote $n_{h}=|\mathcal{V}_{h}|$ and $n_{b}=|\mathcal{V}_{b}|$ the total number of honest and Byzantine nodes, and $N_{h}(i)$ and $N_{b}(i)$ the number of honest and Byzantine neighbors of node $i$. Each agent owns and iteratively updates a local parameter $x_{i}^{t}$. At time $t$, we denote $X^{t}:=[x_{1}^{t},\ldots,x_{n}^{t}]^{T}\in\mathbb{R}^{\mathcal{V}\times d}$ the parameter matrix, which can be split into honest and Byzantine parameter matrices by writing $X=\begin{pmatrix}X^{t}_{h}\\ X^{t}_{b}\end{pmatrix}$, where $X^{t}_{h}=[(x^{t}_{i})_{i\in\mathcal{V}_{h}}]^{T}$ (resp. $X^{t}_{b}$) is the sub-matrix containing the models held by the honest workers (resp. Byzantine). Note that Byzantine nodes (in opposition to malicious nodes) can declare different values to each of their neighbors. As such, only the number and positions of connections with Byzantine nodes matter, and we can w.l.o.g identify the number of edges linking Byzantine and honest nodes $|\mathcal{E}_{b}|$ to the number of Byzantine workers $n_{b}$.
@@ 2.1 Gossip - average consensus case
In gossip algorithms, information is shared between neighbors by performing local averaging steps. As previously explained, we focus on the average consensus problem, which is an important special case of Equation 1.
Definition 2.1 (Average consensus problem (ACP)).
The ACP consists in computing ${\overline{x^{*}}=\frac{1}{n_{h}}\sum_{i\in\mathcal{V}_{h}}x_{i}^{*}}$, where each node $i\in\mathcal{V}_{h}$ holds a value $x_{i}^{*}$. Equivalently, it consists in solving Problem (1) with $f_{i}(x)=\|x-x_{i}^{*}\|^{2}$.
We define a gossip matrix as follows:
Definition 2.2 (Gossip matrix, spectral gap).
@@
A matrix ${W\in\mathbb{R}_{+}^{n\times n}}$ is a gossip matrix for a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ if: (i) $W$ is symmetric and doubly-stochastic: $W\textbf{1}_{n}=\textbf{1}_{n}$, and $\textbf{1}_{n}^{T}W=\textbf{1}_{n}^{T}$. (ii) $W$ characterizes connections in the network: $W_{ij}\neq 0$ if and only if $i=j$ or ${(i\sim j)}\in\mathcal{E}$.
@@
Such a $W$ has eigenvalues $1=\mu_{1}(W)\geqslant|\mu_{2}(W)|\geqslant\ldots\geqslant|\mu_{n}(W)|$. The spectral gap of $W$ is defined as $\gamma_{W}:=1-|\mu_{2}(W)|$. If the graph $\mathcal{G}$ is connected, then $\gamma_{W}>0$. The spectral gap is the main quantity characterizing the information propagation in the graph.
The standard gossip algorithm writes (Boyd et al., 2006):
$X^{t+1}=WX^{t},$ (2)
In the absence of Byzantine nodes, gossip averaging (Equation 2) ensures linear convergence of the parameters to the average (Boyd et al., 2006), with $X^{0}=(x_{i}^{*})_{i\in\mathcal{V}}$ at a linear rate: $\sum_{i\in\mathcal{V}}\|x_{i}^{t}-\overline{x^{*}}\|^{2}\leqslant(1-\gamma)^{2t}\sum_{i\in\mathcal{V}}\|x_{i}^{0}-\overline{x^{*}}\|^{2}.$
Example 2.3 (Gossip from Graph Laplacian).
@@
Let us consider the Laplacian matrix of the graph $L=D-A$ where $D$ is the diagonal matrix of degrees and $A$ is the adjacency matrix (i.e., $A_{ij}=\mathds{1}_{{(i\sim j)}\in\mathcal{E}}$).
Then, $W_{L}=\operatorname{Id}-\eta L$ is a gossip matrix for $\eta\leqslant\frac{1}{\max_{i}(D_{ii})}$. We then have that $\gamma_{W_{L}}=\eta{\mu_{\min}^{+}(L)}$, with $\mu_{\min}^{+}(L)$ the smallest positive eigenvalue of $L$.
@@
Unfortunately, standard gossip communications use non-robust averaging steps, and so they are vulnerable to a single Byzantine node, which can drive all honest nodes to any position (Blanchard et al., 2017).
2.2 Clipped gossip: gossiping robustly
To remedy this vulnerability problem, He et al. (2022) propose a computationally-efficient solution for robust gossip communication. Their algorithm extends self-centered clipping (Karimireddy et al., 2021) to the decentralized case: nodes project all parameters they receive on a ball centered on their own parameter. Formally, for $\tau\geqslant 0$ and $x\in\mathbb{R}^{d}$,
$\operatorname{Clip}(x;\tau):=\frac{x}{\|x\|_{2}}\min(\|x\|_{2},\tau)$ (3)
is the projection of $x$ onto the $L_{2}$ ball of radius $\tau$. Then, for $\boldsymbol{\tau}^{t}\in\mathbb{R}^{\mathcal{V}_{h}}$, He et al. (2022)’s communication algorithm, called $\operatorname{ClippedGossip}(\boldsymbol{\tau}^{t})$, writes:
$\begin{cases}x_{i}^{t+1}=x_{i}^{t}+\sum_{j\in\mathcal{V}}W_{ij}\operatorname{Clip}(x_{j}^{t}-x_{i}^{t};\boldsymbol{\tau}_{i}^{t})\text{ if }i\in\mathcal{V}_{h}\\ x_{i}^{t+1}=\,*,\text{ i.e., is arbitrary,}\,\text{ if }i\in\mathcal{V}_{b}.\end{cases}$ (ClippedGossip)
In the limit case of $\tau_{i}^{t}=\infty$ for all $t\geqslant 0$ and $i\in\mathcal{V}_{h}$ we recover the standard gossip averaging of Equation 2, which with Byzantines writes: $\begin{pmatrix}X^{t+1}_{h}\\ \ldots\end{pmatrix}=W\begin{pmatrix}X^{t}_{h}\\ X^{t}_{b}\end{pmatrix},$ where we denote by $\ldots$ the fact that we have no control on the models held by Byzantines, thus considered as arbitrary: $X_{b}^{t+1}=*$.
It can be shown that one step of $\operatorname{ClippedGossip}$ ensures robust contraction, which allows it to be used efficiently as a communication primitive for Byzantine-robust decentralized optimization. For instance, He et al. (2022) combine $\operatorname{ClippedGossip}$ with local computation steps (stochastic gradients with momentum) to solve Problem 1 for generic convex smooth functions $f_{i}$ in the decentralized Byzantine setting. Generic analyses also show that $\operatorname{ClippedGossip}$ allows Byzantine-Resilient Decentralized SGD to converge to a neighborhood of the optimum (Wu et al., 2023). More generally, the idea of self-centered clipping is at the heart of several Byzantine-robust methods (Gorbunov et al., 2022).
Remark 2.4 (Oracle clipping rule in He et al. (2022)).
To ensure their algorithm’s convergence, He et al. (2022) rely on an oracle clipping threshold, written for node $i\in\mathcal{V}_{h}$ as $\boldsymbol{\tau}_{i}^{t}:=\sqrt{\frac{1}{\delta(i)}\sum_{j\in\mathcal{V}_{h}}W_{ij}\mathbb{E}\|x_{i}^{t+1/2}-x_{j}^{t+1/2}\|^{2}_{2}},$ where $\delta(i)=\sum_{j\in\mathcal{V}_{b}}W_{ij}$ is the weight associated to Byzantine neighbors of node $i$ in the gossip matrix $W$, and $x_{i}^{t+1/2}$ is the local model after the optimization step. Not only does each node need to know the variance of the noise of its neighbors, but it also requires to know which nodes are honest. This breaks the fundamental assumption of not knowing the identity (honest or Byzantine) of the nodes. Even if it is only required for setting the clipping threshold, this is still a major problem since the clipping threshold is a very sensitive parameter, and it is unclear what happens, specifically in case of under-clipping. Experiments are performed using a rule of thumb for the clipping threshold, not supported by theory. In contrast, our approach leads to a computable clipping rule which is supported by theory.
2.3 Dual approach
As previously mentioned, the dual approach is very popular in decentralized learning (Jakovetić et al., 2020; Uribe et al., 2020), and is at the heart of key advances, such as acceleration (Scaman et al., 2017; Hendrikx et al., 2019; Kovalev et al., 2020) or variance reduction (Hendrikx et al., 2021). In this section, we show that $\operatorname{ClippedGossip}$ (He et al., 2022) can be understood as Clipped Gradient Descent in the dual, which we will then leverage for the analysis.

### v2

@@ Abstract
Distributed learning has many computational benefits but is vulnerable to attacks from a subset of devices transmitting incorrect information. This paper investigates Byzantine-resilient algorithms in a decentralized setting, where devices communicate directly in a peer-to-peer manner within a communication network. We leverage the so-called dual approach for decentralized optimization and propose a Byzantine-robust algorithm. We provide convergence guarantees in the average consensus subcase, discuss the potential of the dual approach beyond this subcase, and re-interpret existing algorithms using the dual framework. Lastly, we experimentally show the soundness of our method.
@@ 1 Introduction
Distributed optimization algorithms allow to harnessing the multiplication of devices to tackle the increasing computational complexity of machine learning models. Yet, the involvement of many parties in the learning process opens up the threat of misbehaving devices, that can purposely or inadvertently send arbitrarily harmful messages. Such device failures are commonly termed Byzantine (Lamport et al., 1982). This paper studies Byzantine robustness in the decentralized communication paradigm, which alleviates the risk of a single point of failure due to a central server by making devices directly communicate with each other in a peer-to-peer manner.
We consider devices (a.k.a nodes) that communicate synchronously within a communication network, abstracted as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where each vertex is a node that communicates with its neighbors in the graph through the edges. We denote as $\mathcal{V}_{h}$ the Byzantine nodes, and $\mathcal{V}_{h}=\mathcal{V}\backslash\mathcal{V}_{b}$ the remaining nodes, called honest. Similarly, we split the edges in $\mathcal{E}=\mathcal{E}_{h}\cup\mathcal{E}_{b}$, where $\mathcal{E}_{h}$ are the edges linking honest nodes, while edges in $\mathcal{E}_{b}$ link honest to Byzantine nodes. The subgraph of honest nodes $\mathcal{G}_{h}:=(\mathcal{V}_{h},\mathcal{E}_{h})$ is always assumed to connected. For any $i\in\mathcal{V}_{h}$, we denote $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ the local objective function on node $i$, and we study the optimization problem:
@@
We study decentralized algorithms to solve this problem, which rely on the gossip communication protocol (Boyd et al., 2006; Nedic & Ozdaglar, 2009; Duchi et al., 2011; Scaman et al., 2017), in which each node $i\in\mathcal{V}$ maintains a local parameter $x_{i}\in\mathbb{R}^{d}$, and updates this parameter by performing approximate average of this parameter with those of its neighbors in the graph.
@@
Such gossip algorithms naturally arise when solving a well-chosen dual formulation of Equation 1 with standard first-order methods. This dual approach gives a principled framework to design and analyze efficient decentralized algorithms, leading for instance to acceleration (Scaman et al., 2017; Kovalev et al., 2020), variance reduction (Hendrikx et al., 2019) and has strong links with gradient tracking (Jakovetić, 2018). This paper investigates the benefits of this approach in the Byzantine robust setting.
Related works. Since its introduction by Blanchard et al. (2017), the setting of Byzantine-robust learning has been extensively studied. The large majority of works focus on the federated case, in which a central server organizes the training and communicates directly with each node  (Yin et al., 2018; Chen et al., 2017; Alistarh et al., 2018; Li et al., 2019; El-Mhamdi et al., 2020; Karimireddy et al., 2023; 2021; Farhadkhani et al., 2022; Allouah et al., 2023). Some work investigated the robust decentralized optimization setting, either with fully connected networks (El-Mhamdi et al., 2021; Farhadkhani et al., 2023) or with more generic networks (He et al., 2023; Wu et al., 2023; Gaucher et al., 2025). However, those works only skimmed over the challenges of decentralized optimization in the Byzantine setting, and many key optimization techniques such as gradient tracking or acceleration are still unexplored in the presence of adversaries. Developing the dual framework in the presence of adversaries would thus open the way to tackle these challenges.
Contributions. In this paper, we (i) leverage the dual approach to derive $\operatorname{\operatorname{EdgeClippedGossip}}$, a decentralized robust algorithm based on clipped dual gradient descent, (ii) derive clipping rules and their associated convergence guarantees in the average consensus subproblem, in which $\operatorname{\operatorname{EdgeClippedGossip}}$ recovers existing algorithms. Finally, (iii) we highlight the dynamics of the developed method compared to previous approaches and discuss its generalization beyond the average consensus case.
@@
2 Deriving EdgeClippedGossip
This section describes a dual approach to Byzantine robustness. We first introduce a dual gossip algorithm, then show how it can be made robust using $\operatorname{\operatorname{EdgeClippedGossip}}$.
Setting: We denote $n_{h}=|\mathcal{V}_{h}|$ and $n_{b}=|\mathcal{V}_{b}|$ the total number of honest and Byzantine nodes, and $N_{h}(i)$ and $N_{b}(i)$ the number of honest and Byzantine neighbors of node $i$. Each agent owns and iteratively updates a local parameter $x_{i}^{t}$. At time $t$, we denote $X^{t}:=[x_{1}^{t},\ldots,x_{n}^{t}]^{T}\in\mathbb{R}^{\mathcal{V}\times d}$ the parameter matrix, which we split into honest and Byzantine as $X^{t}=\begin{pmatrix}X^{t}_{h}\\ X^{t}_{b}\end{pmatrix}$, where $X^{t}_{h}=[(x^{t}_{i})_{i\in\mathcal{V}_{h}}]^{T}$ (resp. $X^{t}_{b}$) is the sub-matrix containing the models held by the honest workers (resp. Byzantine). Note that Byzantine nodes can declare different values to each of their neighbors. As such, only the number and positions of edges with Byzantine nodes matter, and we can w.l.o.g identify the number of edges linking Byzantine and honest nodes $|\mathcal{E}_{b}|$ to the number of Byzantine workers $n_{b}$.
@@ 2.1 Gossip - average consensus case
Gossip communications (Boyd et al., 2006) consist in sharing information between neighbors through local averaging steps. The standard gossip protocol writes $X^{t+1}=WX^{t}$, where $W\in[0,1]^{n\times n}$ is a gossip matrix.
Definition 1 (Gossip matrix).
@@
A matrix ${W\in\mathbb{R}_{+}^{n\times n}}$ is a gossip matrix for a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ if: (i) $W$ is symmetric and doubly-stochastic: $W\textbf{1}_{n}=\textbf{1}_{n}$, and $\textbf{1}_{n}^{T}W=\textbf{1}_{n}^{T}$. (ii) $W$ is supported on the edges of the network: $W_{ij}\neq 0$ if and only if $i=j$ or ${(i\sim j)}\in\mathcal{E}$.
@@
Such a matrix $W$ has eigenvalues $1=\mu_{1}(W)\geqslant|\mu_{2}(W)|\geqslant\ldots\geqslant|\mu_{n}(W)|$. The propagation of information in the graph is characterized by the spectral gap of $W$, which is defined as $\gamma_{W}:=1-|\mu_{2}(W)|$. Note that, if the graph $\mathcal{G}$ is connected, then $\gamma_{W}>0$.
Example 1 (Gossip from Graph Laplacian).
@@
Consider the Laplacian matrix of the graph $L=D-A$ where $D$ is the diagonal matrix of degrees and $A$ is the adjacency matrix (i.e., $A_{ij}=\operatorname{1}_{{(i\sim j)}\in\mathcal{E}}$). Then, $W_{L}=\operatorname{Id}-\eta L$ is a gossip matrix for $\eta\leqslant\nicefrac{{1}}{{\mu_{n}(W)}}$, with spectral gap $\gamma_{W_{L}}=\eta{\mu_{\min}^{+}(L)}$, where $\mu_{\min}^{+}(L)$ is the second smallest eigenvalue of $L$.
The average consensus problem is an important special case of Equation 1.
Definition 2.
The approximate average consensus problem (ACP) consists in computing ${\overline{x^{*}}=\frac{1}{n_{h}}\sum_{i\in\mathcal{V}_{h}}x_{i}^{*}}$, where each node $i\in\mathcal{V}_{h}$ holds a value $x_{i}^{*}$, which is equivalent to Problem (1) with $f_{i}(x)=\|x-x_{i}^{*}\|^{2}$. In the presence of Byzantines, only an approximate estimation of $\overline{x^{*}}$ can be found.
@@
When there are no Byzantine nodes, standard gossip-averaging results ensure linear convergence of the parameters to the average (Boyd et al., 2006): $\sum_{i\in\mathcal{V}}\|x_{i}^{t}-\overline{x^{*}}\|^{2}\leqslant(1-\gamma)^{2t}\sum_{i\in\mathcal{V}}\|x_{i}^{0}-\overline{x^{*}}\|^{2}$. Yet, this result is not robust to adversarial nodes: as the standard gossip communication uses averaging steps, it is vulnerable to a single Byzantine node, which can drive all honest nodes to any position (Blanchard et al., 2017).
2.2 Dual approach in decentralized optimization
The use of a dual formulation to solve decentralized optimization is very popular (Jakovetić et al., 2020; Uribe et al., 2020). Not only has it been at the heart of key advances such as acceleration (Scaman et al., 2017; Hendrikx et al., 2019; Kovalev et al., 2020) and variance reduction (Hendrikx et al., 2021), but it also allows us to understand methods such as gradient tracking (Jakovetić, 2018). It is thus natural to investigate whether the dual approach can be used for Byzantine robustness.
We first introduce the dual approach in a Byzantine-free setting. Then, we interpret dual variables and explain their use in the Byzantine-robust setting. Eventually, we instantiate the algorithm in the ACP setting and show that it recovers existing algorithms.